# Eka-Eval [Uncaptioned image] : An Evaluation Framework for Low-Resource Multilingual Large Language Models

### v4

@@ Abstract
The rapid evolution of Large Language Models’ has underscored the need for evaluation frameworks that are globally applicable, flexible, and modular, and that support a wide range of tasks, model types, and linguistic settings. We introduce Eka-Eval, a unified, end-to-end framework that combines a zero-code web interface and an interactive CLI to ensure broad accessibility. It integrates 50+ multilingual benchmarks across nine evaluation categories, supports local and proprietary models, and provides 11 core capabilities through a modular, plug-and-play architecture. Designed for scalable, multilingual evaluation with support for low-resource multilingual languages, Eka-Eval is, to the best of our knowledge, the first suite to offer comprehensive coverage in a single platform. Comparisons against five existing baselines indicate improvements of at least 2x better on key usability measures, with the highest user satisfaction, faster setup times, and consistent benchmark reproducibility.
@@ 1 Introduction
Large Language Models’ (LLMs) growing capabilities continue to reshape NLP, enabling impressive generalization across diverse tasks, including instruction following, reasoning, summarization, translation, and tool use. With the advent of general-purpose foundation models such as GPT-4 (achiam2023gpt), Claude (anthropic2024claude35sonnet), Gemini (team2023gemini), Deepseek (liu2024deepseek), and Llama-3 (touvron2023llama), the focus of research has increasingly shifted from building task-specific models to systematically evaluating these powerful systems. Evaluation plays a critical role not only in measuring progress but in identifying capabilities, exposing limitations, and informing deployment strategies.
@@
Framework
Custom
Data
Custom
Models
Custom
Prompt
Quant-
ization
Long
Context
Tool
Use
Dist.
Infer.
Visual
Analysis
Zero-Code
UI
Interact.
CLI
Low-Resource
Multilingual
@@
Table 1: Feature comparison with state-of-the-art frameworks. Takeaway: Eka-Eval seamlessly integrates an accessible dual interface (Zero-Code UI + Interactive CLI) and cutting-edge features (Long Context, Tool Use) with Low-Resource Multilingual benchmarks.
An Alarming Need: Existing evaluation frameworks for LMs, such as HELM (liang2022helm), EleutherAI Harness (gao2021framework), lm-eval-harness (gao2021framework), and OpenCompass (wu2023opencompass), predominantly target high-resource languages and rely on command-line interfaces with limited benchmark coverage (watts2024pariksha). These frameworks create significant barriers for non-technical users who require graphical interfaces and researchers working with low-resource and multilingual settings. The challenge is especially pronounced in linguistically diverse regions globally, where numerous languages serve large populations; yet, existing evaluation frameworks lack comprehensive multilingual support and culturally grounded benchmarks, limiting rigorous assessment of language model capabilities across diverse linguistic settings (pomerenke2025ai; xuan-etal-2025-mmlu; vayani2025all).
Motivation: Existing evaluation frameworks suffer from several limitations such as (1) absence of user-friendly interfaces, requiring extensive command-line expertise, broken-pipelines and manual configuration (gao2021framework; liang2022holistic), (2) restricted flexibility in benchmark selection and task customization (wang2018glue; srivastava2022beyond), (3) insufficient task diversity, with fragmented support for advanced capabilities like long-context reasoning and tool use (qin2023toolllm; zhang2024bench), and (4) limited multilingual coverage, particularly for low-resource Indian languages (hu2020xtreme; ahuja2023mega).
@@
  • We propose Eka-Eval, a unified and modular evaluation framework that integrates over 50 multilingual benchmarks across nine major evaluation categories.
  • We introduce a first-ever Zero-Code UI for benchmark evaluation in diverse and low-resource multilingual settings, along with visualisations, leaderboard, and an interactive CLI, making large-scale evaluation accessible to both non-technical and technical users.
@@ 2 Relevant Work
General-Purpose Frameworks: Early multi-task suites such as GLUE (wang2018glue) and SuperGLUE (wang2019superglue) established foundational standards for evaluation. More comprehensive frameworks, including HELM (liang2022helm), BIG-Bench (srivastava2022beyond), lm-eval-harness (gao2021framework), and OpenAI Evals (openai2023evals) broaden task coverage and introduce extensible evaluation pipelines. yet typically require substantial effort and provide limited multilingual capabilities. Modular ecosystems such as OpenCompass (wu2023opencompass) and FreeEval (yu2024freeeval) offer distributed inference and advanced features, but still lack unified evaluation and zero-code accessibility.
Specialized Capability Benchmarks: Several domain-specific benchmarks target advanced LLM abilities, such as tool use in ToolBench (qin2023toolbench), agentic behavior in API-Bank (li2023api) and InfiniteBench (zhang2024bench), long-context reasoning in RULER (hsieh2024ruler), and mathematical and coding competencies in GSM8K (cobbe2021training) and HumanEval (chen2021evaluating). While powerful, these benchmarks are fragmented across independent repositories, requiring users to mainitain multiple codebases to conduct a thorough evaluation.
Multilingual Evaluation: Multilingual suites such as XTREME (hu2020xtreme) and MEGA (ahuja2023mega) cover 30+ languages but often rely on machine-translated content that lacks cultural grounding. Recent efforts of MCEval (Huang2025MCEvalAD) (13 cultures, 39K dynamic instances), GlotEval (Luo2025GlotEvalAT) (language-agnostic, culturally adaptive prompts), BLEnD (myung2024blend) (16 cultures, 52K native QA pairs), CulturalBench (chiu2024culturalbench) (45 global regions), and ALM-Bench (vayani2025all) (100 languages with cultural images) address this gap. However, these benchmarks still face insufficient data for low-resource languages, cultural biases favor dominant dialects, and evaluation metrics struggle to capture cultural appropriateness.
Eka-Eval mitigates these gaps by unifying 50+ diverse benchmarks within a single modular workflow. It offers dual accessibility (zero-code UI + interactive CLI), advanced capabilities (long-context, tool-use, distributed inference), and multilingual language support for comprehensive evaluation.
@@ B. Advanced Capabilities:
(8) Visual Analysis: Generation of interpretable visualizations (bar charts, radar plots, heatmaps, etc.).
@@ C. Usability & Specialization:
(9) Zero-Code UI: Web-based interface enabling evaluation without programming.
@@
Table 1 presents a feature-level comparison demonstrating that Eka-Eval is the only framework offering complete coverage across technical, linguistic, and usability dimensions. While lm-eval-harness, OpenCompass, and HELM provide strong technical capabilities, they require substantial expertise and suffer from usability challenges. indic-eval, though tailored to Indic languages, supports only seven benchmarks and lacks broader functionality. Critically, existing frameworks lack zero-code web interfaces, restricting accessibility to users with programming expertise. Eka-Eval uniquely combines comprehensive evaluation capabilities with an intuitive zero-code UI, significantly lowering barriers to rigorous multilingual assessment.
@@ 4.1 Design and Implementation
Eka-Eval is architected as a modular, extensible evaluation framework that balances practical usability with comprehensive benchmark coverage. It provides a zero-code, web-based interface that allows users to run benchmarks, adjust parameters, and visualize results through interactive analytics and leaderboards, while a CLI enables detailed configuration and large-scale evaluations. The framework supports flexible model selection for both local and API-based models and is built around three core principles: modularity for seamless extension and customization; accessibility to support diverse environments, including low-resource language settings; and comprehensiveness to cover a wide range of capabilities and multilingual benchmarks, with particular attention to underserved areas such as long-context reasoning and tool use.
@@ 4.2 System Architecture
The framework has a layered architecture with four main components (illustrated in Figure 1):
@@ 4.2.3 Model Interface Layer
Table 2: Average participant ratings of the evaluation frameworks by eleven participants (mean $\pm$ standard deviation; Likert scale: 1–5, 1 = poor and 5 = excellent). Takeaway: Eka-Eval achieves the highest ratings across all functionalities, with participants consistently perceiving it as superior across the six assessed capabilities.
@@
Table 3: Comparison of time taken for installation and configuration across six frameworks reported across five successful runs on GPU. Takeaway: Eka-Eval records the lowest setup time and exhibits consistently low variance.
@@ 4.2.4 Results Processing System
Metrics Calculator: Employs HuggingFace’s industry-standard evaluate library to compute essential metrics such as accuracy, BLEU, F1-score, exact match, and Pass@1, ensuring consistency with widely adopted benchmarks and standard evaluation practices across diverse LLM tasks.
@@ 4.2.5 Interface and Deployment Layer
Full-Stack Architecture: A decoupled React frontend and FastAPI backend, served through Nginx, provide a stable and scalable deployment layer.
Real-Time Telemetry: WebSocket-based streaming delivers live inference logs and GPU status directly to the UI, eliminating the need for CLI monitoring.
@@
Interactive Dashboard: A “Zero-Code” interface with visual comparisons and leaderboards enables users to run and analyze evaluations without technical overhead.
@@ 4.3 Comprehensive Benchmark Coverage
Eka-Eval covers nine major evaluation categories with comprehensive benchmark support across 50+ benchmarks (See Appendix A).
@@
    Code Generation and Programming: Programming abilities are assessed using HumanEval (chen2021evaluating), MBPP (austin2021program), HumanEval+ (liu2023your), and MBPP EvalPlus (liu2023your) with Pass@1 metrics.
@@
    Reading Comprehension: Text understanding is evaluated using SQuAD (rajpurkar2018know), QuAC (choi2018quac), and BoolQ (clark2019boolq) with F1 and exact match metrics.
@@
    Long-Context Understanding: For extended context, we include ZeroSCROLLS (shaham2023zeroscrolls), Needle-in-a-Haystack (wang2024multimodal), and InfiniteBench (zhang2024bench).
@@
    General Reasoning: Foundational capabilities are tested via MMLU (hendrycks2021mmlu), MMLU-Pro (wang2024mmlu), IFEval (zhou2023instruction), BBH (suzgun2022challenging), and AGI-Eval (zhong2023agieval).
@@
    Multilingual Language Support: Eka-Eval features the largest unified Multilingual suite (20 benchmarks), including: (i) Knowledge: IndicMMLU-Pro (sankalp2025indicmmlupro), MMLU-IN (hendrycks2020measuring), TriviaQA-IN (joshi2017triviaqa), MILU (verma2024milu). (ii) Reasoning: HellaSwag-IN (zellers2019hellaswag), ARC-C IN (clark2018think), IndicCOPA (kakwani2020indicnlpsuite), GSM8K-IN (cobbe2021training). (iii) Reading & QA: BoolQ-IN (clark2019boolq), XQuAD-IN (artetxe2019cross), XorQA-IN (asai2021xor), Indic-QA (singh2025indicqabenchmarkmultilingual). (iv) Generation (NLG): Flores-IN (goyal-etal-2022-flores), IndicParaphrase, IndicWikiBio, IndicQuestionGeneration, IndicSentenceSummarization, IndicHeadlineGeneration (singh2024indicgenbench). (v) NLU: IndicNER, IndicSentiment, IndicGLUE (kakwani2020indicnlpsuite).
   Framework    Winogrande    PIQA    ARC-IN    MMLU-IN
   lm-eval-harness    65.8 $\pm$ 6.0    74.8 $\pm$ 3.5    –    –
   OpenCompass    61.1 $\pm$ 5.0    73.3 $\pm$ 4.3    –    –
   indic-eval    65.8 $\pm$ 1.9    78.7 $\pm$ 1.4    28.1 $\pm$ 2.3    35.0 $\pm$ 2.1
   FreeEval    60.0 $\pm$ 4.5    54.5 $\pm$ 3.2    –    –
   Eka-Eval    66.5 $\pm$ 2.5    77.5 $\pm$ 1.5    31.5 $\pm$ 3.5    33.5 $\pm$ 0.5
4.4 Extensibility and Customization
Eka-Eval emphasizes extensibility through a low-code plugin architecture that enables easy integration of new benchmarks and custom metrics with minimal modifications, managed via a hierarchical JSON configuration system that supports complex setups such as parameter sweeps and prompt variations. The framework provides a zero-code UI that enables real-time, interactive adjustment of inference parameters (e.g., temperature, batch size) and evaluation strategies, eliminating the need for users to manage configuration files directly.
@@ 5 Experiments
To assess the effectiveness and usability of Eka-Eval, we conducted a comprehensive evaluation combining benchmark coverage analysis and user-centered feedback across six prominent frameworks: lm-eval-harness, OpenCompass, HELM, FreeEval, indic-eval, and Eka-Eval.
@@
Eleven graduate-level computer science researchers with demonstrated proficiency in Git/GitHub evaluated each framework in a single-blind setup. Participants followed a standardized evaluation protocol in a controlled environment. They installed each framework, integrated HuggingFace models - including gemma-2-2b (team2024gemma) and sarvam-1B sarvamai_sarvam1_2024, and executed the recommended workflows on four diverse benchmarks: WinoGrande, PIQA, ARC-C-IN and MMLU-IN. They then rated the framework on five criteria using a Likert scale ranging from 1 to 5:
EKA-EVAL Analysis: All ratings were recorded to enable direct comparison across frameworks. Detailed instructions and a standardized rating form were provided to ensure consistency and reproducibility, with explicit dos and don’ts specified for all participants. They were permitted to consult official documentation but were prohibited from seeking external technical support beyond publicly available framework resources.
Participants assessed each framework across six evaluation criteria:
@@
    Extensibility: Ease of customizing the framework to add new prompt templates, models, datasets, or evaluation metrics.
@@
    Multilingual Language Support: Support for multilingual benchmarks such as ARC-IN, FLORES, and similar. Evaluators assessed the availability of prebuilt resources.
Table 2 presents average participant ratings across six evaluation criteria using a 1-5 Likert scale (1= poor or absence of functionality, 5 =excellent or flexible capability). Eka-Eval achieved the highest ratings across all criteria, establishing superior usability and capability. Moreover, Table 3 summarizes the average setup times across six frameworks, which varied widely due to dependency issues and unstable codebases in several baselines. Nine of the eleven participants reported that HELM suffered from poor documentation and a broken pipeline, while OpenCompass required manual code edits and indic-eval experienced installation difficulties. Eka-Eval achieved the shortest setup time. Finally, Table 4 presents reproduced benchmark results on gemma-2-2b, showing how they align with broader framework differences and highlighting Eka-Eval’s consistency across global-multilingual benchmarks. HELM was excluded from this comparison, as it neither supports these benchmarks nor produced functional results despite repeated attempts.

### v5

@@ Abstract
The rapid evolution of Large Language Models has underscored the need for evaluation frameworks that are globally applicable, flexible, and modular, and that support a wide range of tasks, model types, and linguistic settings. We introduce Eka-Eval, a unified, end-to-end framework that combines a zero-code web interface and an interactive CLI to ensure broad accessibility. It integrates 55+ diverse benchmarks across nine evaluation categories, supports local and proprietary models, and provides 11 core capabilities through a modular, plug-and-play architecture. Designed for scalable, multilingual evaluation with support for low-resource multilingual languages, Eka-Eval is, to the best of our knowledge, the first suite to offer comprehensive coverage in a single platform. Comparisons against five existing baselines indicate improvements of at least 2x better on key usability measures, with the highest user satisfaction, faster setup times, and consistent benchmark reproducibility.
@@ 1 Introduction
Large Language Models (LLMs) growing capabilities continue to reshape NLP, enabling impressive generalization across diverse tasks, including instruction following, reasoning, summarization, translation, and tool use. With the advent of general-purpose foundation models such as GPT-4 (achiam2023gpt), Claude (anthropic2024claude35sonnet), Gemini (team2023gemini), Deepseek (liu2024deepseek), and Llama-3 (touvron2023llama), the focus of research has increasingly shifted from building task-specific models to systematically evaluating these powerful systems. Evaluation plays a critical role not only in measuring progress but in identifying capabilities, exposing limitations, and informing deployment strategies.
@@
Framework Custom Data Custom Models Custom Prompt Quant- ization Long Context Tool Use Dist. Infer. Visual Analysis Zero-Code UI Interact. CLI Low-Resource Multilingual
@@
Table 1: Feature comparison with state-of-the-art frameworks. Takeaway: Eka-Eval is the only framework to achieve full coverage across all eleven dimensions, uniquely combining a zero-code UI and low-resource multilingual support with advanced capabilities such as long-context and tool-use evaluation.
An Alarming Need: Existing evaluation frameworks for LMs, such as HELM (liang2022holistic), lm-eval-harness (gao2021framework), OpenCompass (wu2023opencompass), DeepEval (deepeval), and UltraEval (he-etal-2024-ultraeval) predominantly target high-resource languages and rely on command-line interfaces with limited benchmark coverage (watts2024pariksha). These frameworks create significant barriers for non-technical users who require graphical interfaces and researchers working with low-resource and multilingual settings. The challenge is especially pronounced in linguistically diverse regions globally, where numerous languages serve large populations; yet, existing evaluation frameworks lack comprehensive multilingual support and culturally grounded benchmarks, limiting rigorous assessment of language model capabilities across diverse linguistic settings (pomerenke2025ai; xuan-etal-2025-mmlu; vayani2025all).
Motivation: Existing evaluation frameworks suffer from several limitations, such as (1) absence of user-friendly interfaces, requiring extensive command-line expertise, broken-pipelines, and manual configuration (gao2021framework; liang2022holistic), (2) restricted flexibility in benchmark selection and task customization (wang2018glue; srivastava2022beyond), (3) insufficient task diversity, with fragmented support for advanced capabilities like long-context reasoning and tool use (qin2023toolbench; zhang-etal-2024-bench), and (4) limited multilingual coverage, particularly for low-resource languages (ahuja-etal-2023-mega; luo-etal-2025-gloteval).
@@
  • We propose Eka-Eval, a unified and modular evaluation framework that integrates over 55 multilingual benchmarks across nine major evaluation categories.
  • We introduce an open-source LLM evaluation framework that unifies a zero-code UI, an interactive CLI, and built-in support for low-resource multilingual settings, along with visualisations, and leaderboard, making large-scale evaluation accessible to both non-technical and technical users.
@@ 2 Relevant Work
General-Purpose Frameworks: Early multi-task suites such as GLUE (wang2018glue) and SuperGLUE (wang2019superglue) established foundational standards. More comprehensive frameworks, including HELM (liang2022holistic), BIG-Bench (srivastava2022beyond), and lm-eval-harness (gao2021framework), introduced extensible evaluation pipelines but typically require substantial technical effort and provide limited multilingual capabilities. Modular ecosystems such as FreeEval (yu2024freeeval) and OpenCompass (wu2023opencompass) offer advanced features like distributed inference. While OpenCompass covers extensive benchmarks and zero-code UI for limited models, custom evaluations still require CLI interaction or manual YAML configuration.
Specialized Capability Benchmarks: Several domain-specific benchmarks target advanced LLM abilities, such as tool use in ToolBench (qin2023toolbench), agentic behavior in API-Bank (li2023api) and InfiniteBench (zhang-etal-2024-bench), long-context reasoning in RULER (hsieh2024ruler), and mathematical and coding competencies in GSM8K (cobbe2021training) and HumanEval (chen2021evaluating). While powerful, these benchmarks are fragmented across independent repositories, requiring users to maintain multiple codebases to conduct a thorough evaluation.
Multilingual Evaluation: Multilingual suites such as XTREME (hu2020xtreme) and MEGA (ahuja-etal-2023-mega) cover 30+ languages but often rely on machine-translated content that lacks cultural grounding. Recent efforts of MCEval (Huang2025MCEvalAD) (13 cultures, 39K dynamic instances), GlotEval (luo-etal-2025-gloteval) (language-agnostic, culturally adaptive prompts), BLEnD (myung2024blend) (16 cultures, 52K native QA pairs), CulturalBench (chiu2024culturalbench) (45 global regions), and ALM-Bench (vayani2025all) (100 languages with cultural images) address this gap. However, these benchmarks still face insufficient data for low-resource languages, cultural biases favor dominant dialects, and evaluation metrics struggle to capture cultural appropriateness.
Existing frameworks are often complex and poorly modularized, hindering seamless incorporation into research workflows (he-etal-2024-ultraeval). Furthermore, most toolkits also remain disproportionately focused on English and a handful of high-resource languages, making non-English-centric evaluations less practical for many researchers (luo-etal-2025-gloteval), and none offer a zero-code UI or native support for low-resource languages, creating a gap for practitioners evaluating models in low-resource settings without dedicated systems expertise.
Eka-Eval addresses these gaps by unifying 55+ benchmarks across nine evaluation categories within a single modular workflow, with a strong focus on low-resource multilingual evaluation. It is, to our knowledge, the only framework to pair a zero-code web UI along with interactive CLI, easy-to-use framework making rigorous evaluation accessible to non-technical users.
@@ B. Advanced Capabilities:
(8) Visual Analysis: Generation of interpretable visualizations (bar charts, plots, heatmaps, etc.).
@@ C. Usability & Specialization:
(9) Zero-Code UI: Web-based interface enabling end-to-end evaluation workflows without coding.
@@
Table 1 demonstrates that Eka-Eval achieves complete technical, linguistic, and usability coverage. While lm-eval-harness and HELM offer strong capabilities, they demand substantial programming expertise. OpenCompass includes a zero-code UI, but restricts it to leaderboard models and pre-configured benchmarks. OpenAI Evals and DeepEval support custom pipelines but lack quantization, long-context, and multilingual capabilities. Similarly, UltraEval and GlotEval provide provide robust customization and visual analysis (with the latter targeting low-resource languages) but but lack tool-use, long-context evaluation, and a zero-code UI. indic-eval is limited to seven benchmarks and offers no general-purpose capabilities. Critically, no existing framework pairs a zero-code UI with native low-resource multilingual support; Eka-Eval addresses both, lowering barriers to rigorous multilingual assessment.
@@ 4.1 Design and Implementation
Eka-Eval is architected as a modular, extensible evaluation framework that balances practical usability with comprehensive benchmark coverage. It provides a zero-code, web-based interface that allows users to run benchmarks, adjust parameters, and visualize results through interactive analytics and leaderboards, while a CLI enables detailed configuration and large-scale evaluations. The framework supports flexible model selection for both local and API-based models and is built around three core principles: modularity for seamless extension and customization; accessibility to support diverse environments, including low-resource language settings; and comprehensiveness to cover a wide range of capabilities and low-resource multilingual benchmarks, with particular attention to underserved areas such as long-context reasoning and tool use.
@@ 4.2 System Architecture
The framework has a layered architecture with four main components (as illustrated in Figure 1):
@@ 4.2.3 Model Interface Layer
Table 2: Average participant ratings of the evaluation frameworks by eleven participants (mean $\pm$ standard deviation; Likert scale: 1–5, 1 = poor and 5 = excellent). Takeaway: Eka-Eval achieves the highest mean ratings across all six dimensions, with notably stronger scores on Zero-Code UI and Global Low-Resource Language Support.
@@
Table 3: Comparison of time taken for installation and configuration across six frameworks reported across five successful runs on GPU-setup. Takeaway: Eka-Eval records the lowest setup time and variance.
@@ 4.2.4 Results Processing System
Metrics Calculator: Employs HuggingFace’s industry-standard evaluate library to compute essential metrics such as accuracy, BLEU (papineni2002bleu), F1-score (10.1145/3606367), exact match (rajpurkar2016squad), and Pass@1 (chen2021evaluating), ensuring consistency with widely adopted benchmarks and standard evaluation practices across diverse LLM tasks.
@@ 4.2.5 Interface and Deployment Layer
Full-Stack Architecture: A decoupled React fedosejev2015react frontend and FastAPI tiangolo2018fastapi backend, served through Nginx, provide a stable and scalable deployment layer.
Real-Time Telemetry: WebSocket-based streaming delivers live inference logs and GPU status directly to the UI, eliminating the need for CLI.
@@
Interactive Dashboard: A “Zero-Code” user-interface with visual comparisons and live leaderboards enables users to run and analyze evaluations without technical expertise.
@@ 4.3 Comprehensive Benchmark Coverage
Eka-Eval covers nine major evaluation categories with comprehensive benchmark support across 55+ benchmarks (See Appendix A).
@@
    Code Generation and Programming: These are assessed using HumanEval (chen2021evaluating), MBPP (austin2021program), HumanEval+ (liu2023your), PythonSaga (yadav-etal-2024-pythonsaga) and MBPP EvalPlus (liu2023your) with Pass@1 metrics.
@@
    Reading Comprehension: Text understanding is evaluated using SQuAD (rajpurkar2018know), QuAC (choi-etal-2018-quac), and BoolQ (clark2019boolq) with F1 and exact match metrics.
@@
    Long-Context Understanding: For extended context, we include ZeroSCROLLS (shaham2023zeroscrolls), Needle-in-a-Haystack (wang-etal-2025-multimodal), and InfiniteBench (zhang-etal-2024-bench).
@@
    General Reasoning: Foundational capabilities are tested via MMLU (hendrycks2021mmlu), MMLU-Pro (wang2024mmlu), IFEval (zhou2023instruction), BBH (suzgun2022challenging), and AGI-Eval (zhong-etal-2024-agieval).
@@
    Multilingual and Low-Resource Language Support: LLMs are known to underperform on languages spoken across South Asia, Africa, and Southeast Asia, yet no existing evaluation framework provides sufficient coverage to diagnose this gap (luo-etal-2025-gloteval; kakwani2020indicnlpsuite). Eka-Eval features the largest unified Multilingual suite (23 benchmarks), including: (i) Knowledge: IndicMMLU-Pro (sankalp2025indicmmlupro), MMLU-IN (hendrycks2021measuring), TriviaQA-IN (joshi2017triviaqa), MILU (verma-etal-2025-milu) spanning Kannada, Odia, Malayalam, and Urdu. (ii) Reasoning: HellaSwag-IN (zellers2019hellaswag), ARC-C-IN (clark2018think), IndicCOPA (kakwani2020indicnlpsuite), XCOPA (ponti-etal-2020-xcopa) in Quechua, Indonesian, and Swahili, and GSM8K-IN (cobbe2021training) for reasoning across Indic languages. (iii) Reading & QA: Belebele (bandarkar-etal-2024-belebele) covering 122 languages including Yoruba, Oromo, and Indonesian; BoolQ-IN (clark2019boolq), XQuAD-IN (artetxe2020cross) in Hindi and Greek, XorQA-IN (asai2021xor) in Bengali and Telugu, and Indic-QA (singh2025indicqabenchmarkmultilingual). (iv) Generation (NLG): Flores-IN (goyal-etal-2022-flores) for translation across 200 languages; IndicParaphrase, IndicWikiBio, IndicQuestionGeneration, IndicSentenceSummarization, and IndicHeadlineGeneration (singh2024indicgenbench) in Odia, Assamese and Punjabi. (v) NLU: IndicNER and IndicSentiment in 11 languages, IndicGLUE (kakwani2020indicnlpsuite), and XNLI (conneau-etal-2018-xnli) for cross-lingual inference in Arabic, Swahili, and Urdu.
@@ 5 Experiments
To assess the effectiveness and usability of Eka-Eval, we conducted a comprehensive evaluation combining benchmark coverage analysis and user-centered feedback across six prominent frameworks: lm-eval-harness, OpenCompass, HELM, FreeEval, indic-eval, and Eka-Eval. These frameworks were selected as representative state-of-the-art baselines spanning across widely adopted general-purpose to specialized multilingual evaluation suites, ensuring a rigorous comparison.
@@
Eleven graduate-level computer science researchers were selected from NLP and software engineering programs. Participants were stratified by prior experience using at least one existing evaluation framework (5 with, 6 without) and evaluated each framework in a single-blind setup. Selection was not tied to knowledge of Eka-Eval, mitigating demand bias; sample size was constrained by available GPU resources. Participants followed a standardized evaluation protocol in a controlled environment. They installed each framework, integrated HuggingFace models - including gemma-2-2b (team2024gemma) and sarvam-1B sarvamai_sarvam1_2024, and executed the recommended workflows on four diverse benchmarks: WinoGrande, PIQA, ARC-C-IN and MMLU-IN.
All ratings were recorded using the standardized protocol (Appendix §D) to ensure direct and consistent comparison across frameworks. Participants assessed each framework across six evaluation criteria and rated them using a Likert scale ranging from 1 to 5:
@@
    Extensibility: Ease of customizing the framework to add new prompt templates, models, benchmarks, or evaluation metrics.
@@
    Multilingual Language Support: Support for multilingual benchmarks such as Belebele, ARC-IN, FLORES, and XNLI.
Table 2 presents average participant ratings across six evaluation criteria using a 1-5 Likert scale (1= poor or absence of functionality, 5= excellent or flexible capability). Eka-Eval achieved the highest ratings across all criteria, establishing superior usability and capability. Moreover, Table 3 summarizes the average setup times across six frameworks, which varied widely due to dependency issues and unstable codebases in several baselines. Nine of the eleven participants reported that HELM suffered from poor documentation and a broken pipeline, while OpenCompass required manual code edits and indic-eval experienced installation difficulties. Eka-Eval achieved the shortest setup time. Table 4 (See Appendix 4 for full results) presents reproduced benchmark scores on gemma-2-2b, showing how they align with broader framework differences and highlighting Eka-Eval’s consistency across global-multilingual benchmarks. HELM was excluded from this comparison, as it neither supports these benchmarks nor produced functional results despite repeated attempts.